### v2

# An Attention-Assisted AI Model for Real-Time Underwater Sound Speed Estimation Leveraging Remote Sensing Sea Surface Temperature Data

Abstract
The estimation of underwater sound velocity distribution serves as a critical basis for facilitating effective underwater communication and precise positioning, given that variations in sound velocity influence the path of signal transmission. Conventional techniques for the direct measurement of sound velocity, as well as methods that involve the inversion of sound velocity utilizing acoustic field data, necessitate on–site data collection. This requirement not only places high demands on device deployment, but also presents challenges in achieving real-time estimation of sound velocity distribution. In order to construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, we propose a self-attention embedded multimodal data fusion convolutional neural network (SA-MDF-CNN) for real-time underwater sound speed profile (SSP) estimation. The proposed model seeks to elucidate the inherent relationship between remote sensing sea surface temperature (SST) data, the primary component characteristics of historical SSPs, and their spatial coordinates. This is achieved by employing CNNs and attention mechanisms to extract local and global correlations from the input data, respectively. The ultimate objective is to facilitate a rapid and precise estimation of sound velocity distribution within a specified task area. The comparative analysis demonstrates that the proposed approach achieves superior performance in terms of both accuracy and stability, exhibiting reduced error rates and enhanced resistance to disturbances when benchmarked against existing advanced techniques.
Index Terms:
self-attention, multimodal data fusion, convolutional neural network (CNN), real-time underwater sound speed profile (SSP) estimation, remote sensing sea surface temperature (SST) data.
I Introduction
Underwater sound velocity construction plays an indispensable role in many applications such as underwater positioning, navigation, timing, communication (PNTC) and target recognition due to its decisive factor in the signal propagation mode [1, 2, 3]. Acoustic propagation speed in marine settings exhibits spatial variability due to the combined effects of temperature, salinity, and pressure [4, 5, 6, 7]. Generally, the sound velocity exhibits a stratified variation with depth, so that sound speed profiles (SSPs) are commonly employed to describe the distribution of sound velocity.
The methodology for acquiring SSPs has been a popular research topic for a long period. In traditional ways, SSPs can be directly measured using various instruments, including the conductivity–temperature–depth profiler (CTD), the expendable CTD (XCTD), and the sound velocity profiler (SVP). Alternatively, SSPs can be derived by inversion based on acoustic field data, which encompasses parameters such as signal transmission time or received signal strength [8, 9, 10, 11, 12]. In general, instrument-based measurements can provide precise SSPs with high–depth resolution, but the measurement process is notably time-intensive. For instance, when the instrument is deployed at a regular speed of 50 meters per minute, it takes at least 80 minutes to measure the sound velocity value within a depth range of 2000 meters to complete the deployment and retrieval of the equipment [11]. As a result, it is difficult to provide real-time sound velocity distributions for underwater communication and positioning through instrument-based measurement. Considering that the distribution of sound velocity can affect the distribution of sound field, researchers have proposed a series of methods to invert the sound velocity distribution by using on-site measured sound field data such as signal propagation time in order to accelerate the process of obtaining sound velocity. There are three main frameworks for inversion methods, including matching field processing (MFP) [13], compressive sensing (CS) [14, 15], and deep learning (DL) [4, 16]. The proposed inversion technique enables fast acquisition of sound speed profiles, offering substantial temporal advantages over conventional conductivity-temperature-depth (CTD) sensors and sound velocity profilers (SVPs) in marine data collection. However, because of the interference of noise in sound field measurement, there is a loss in the accuracy of these sound velocity inversion methods. More importantly, it still takes a certain amount of time to measure the sound field data, and new requirements have been put forward for the deployment of expensive sound field measurement equipment. The implementation of SSP inversion techniques is inherently constrained by their dependence on active sonar measurements, thereby limiting their applicability in regions beyond the operational range of underwater acoustic monitoring networks.
Over the past few years, the fast development of underwater automated sensing equipment has provided increasingly rich vertical observation reference data for exploring the distribution of underwater sound velocity. To achieve fast estimation of SSPs and eliminate the need for underwater on-site operations, various methods have emerged that use empirical sound velocity distribution data or remote sensing data. [17, 18, 19] utilize empirical SSPs data to capture the changes in sound velocity from a time series perspective, enabling real-time estimation and prediction of sound velocity distribution in designated ocean areas. These SSP prediction methods have a single type of input data and low complexity of model construction. However, because of the low time resolution of historical samples, its accuracy performance is not good enough in estimating the sound velocity within a small-scale time range. In fact, the variation of sound velocity distribution in a small-scale spatial area is mainly reflected in the shallow water part, as it is more significantly affected by temperature changes. Marine remote sensing technology provides real-time and reliable high spatiotemporal resolution sea surface temperature (SST) data [20], which can offer initial sea surface condition constraints for estimating underwater sound velocity distribution. [21] proposed a self-organizing map (SOM) neural network model that combines empirical sound velocity data and remote sensing SST data, but the model mainly focuses on the local characteristics of sound velocity distribution, which not only fails to capture the influence of SST data on deep ocean sound velocity, but also fails to capture the spatial correlation of sound velocity distribution. As a result, the accuracy performance has not been significantly improved compared to prediction methods.
In order to fast and accurately estimate the SSP of a given task area without on-site underwater data measurement, we propose an interpretable self-attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model in this paper to deeply capture the intrinsic relationships among empirical SSP data, remote sensing SST data, and spatial coordinates. The core idea is to capture the local feature correlation of regional sound velocity distribution through CNN, and enhance the learning ability of global feature correlation of multimodal data through attention mechanism. To evaluate the effectiveness of the proposed SA-MDF-CNN model, we conducted experiments using Argo data from the Pacific Ocean and measured data from the South China Sea in 2023. Results show that SA-MDF-CNN can achieve lower root mean square error (RMSE) than other spatial SSP construction methods. The contribution of this paper can be summarized as follows:
  • To achieve real-time estimation of SSP without on-site underwater data measurement, we propose an attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model. We take into account the mutual influence among empirical sound velocity distribution, SST, and spatial position, and the attention module is designed to enhance the global feature learning ability of the model.
  • To evaluate the accuracy performance of the model, we conducted experiments using Argo data from the Pacific region and further validated it through sea trials in the South China Sea in 2023. The RMSE of both experimental results is superior to other spatial SSP constructing methods.
  • To enhance the interpretability of the model, the weights of the model parameters are visualized. As the number of training iterations increases, the model parameter weights gradually focus on the shallow water part. This is reasonable because through the empirical orthogonal function (EOF) decomposition of historical SSPs, it can be seen that the differences in sound velocity distribution are more prominent in shallow water. Therefore, the weight visualization results indicate that the model can effectively capture the characteristics of sound velocity distribution.
The structure of this article is arranged as follows. Section 2 provides a detailed introduction to relevant works. The overall architecture and functional module description of the SA-MDF-CNN model are provided in Section 3. Section 4 gives the experimental findings, and the final section, Section 5, presents the concluding remarks.
II Related works
The distribution of underwater sound velocity plays an essential role in the energy distribution and propagation trajectory of acoustic signals. Therefore, real-time and accurate estimation of sound velocity distribution has significant value for applications based on communication and positioning technologies such as underwater PNTC systems, and target recognition systems.
The traditional way of obtaining sound velocity distribution is usually through direct measurement using shipborne CTD or SVP equipment, which has the advantage of high accuracy [22, 23]. However, it comes with high economic costs and measurement time expenses. With the development of sensing and network technology, many kinds of underwater environmental observation systems have been established to enhance human understanding of the ocean. In 1979, MUNK and Wunsch first put forward the concept of ocean acoustic tomography that using acoustic field observation information to invert the sound velocity distribution [24, 25]. Since then, there have been three mainstream frameworks for sound velocity inversion, namely MFP [13], CS [14, 15], and DL [4, 16]. The MFP framework mainly consists of four steps. Firstly, EOF decomposition is adopted to extract the principal component features of the regional sound velocity distribution. Then, different feature combinations are generated to form candidate SSPs. Next, the sound field distribution is simulated based on ray theory. Finally, the simulated sound field is matched with the measured sound field to determine the optimal estimation of SSP. To accelerate the search speed of the optimal candidate SSP, Tolstoy introduced simulated annealing algorithm in the MFP, which improved the algorithm execution efficiency but resulted in suboptimal solutions [13]. Afterwards, other heuristic algorithms were combined with MFP frameworks for SSP inversion, but they also obtained suboptimal solutions [26, 27]. To further improve the efficiency of inversion algorithm execution, Bianco [15] and Choo [14] proposed SSP inversion framework based on CS, respectively. In the CS framework, the mapping relationship from the sound field distribution to the sound velocity distribution is established through a matrix, which eliminates the search process for matching terms. However, the matrix relationship introduces linear approximation, thus sacrificing inversion accuracy.
In recent years, deep learning theory has developed rapidly and achieved significant results in constructing unknown nonlinear relationships between different data [28, 29]. In addition, long-term underwater environmental observations have accumulated a large amount of data for marine hydrological research, laying a data foundation for the application of deep learning underwater. To address the drawbacks of computational efficiency and accuracy loss in MFP and CS frameworks, we proposed a auto-encoding feature mapping neural network model for SSP inversion in our early work [4], in which the auto-encoder was created to extract deep robust features so as to reduce the impact of acoustic field measurement errors on the accuracy of sound velocity inversion. Considering the insufficient accumulation of historical sound velocity data in some ocean areas, the DL model is prone to over-fitting and reduces accuracy performance. Therefore, we further propose a few-shot meta learning framework to accelerate the convergence speed of the model [16]. However, the above-mentioned methods based on MFP, CS, and DL all rely on real-time measured acoustic field data, which puts high demands on the deployment of underwater observation equipment. Therefore, these methods not only face high equipment economic costs, but also have limited application scope for areas without sonar measurement systems.
Figure 1: SSP Estimation Structure based on SA-MDF-CNN.
Nowadays, how to eliminate the need for underwater on-site data measurement has become a research hotspot in the field of SSP inversion. Based on this requirement, Liu et al. and Lu et al. both proposed a long-short term memory (LSTM) neural network model for SSP prediction [17, 18], which only requires empirical SSP data. However, due to the low temporal resolution of prior data in most marine areas, this prediction method can only describe the overall trend of sound velocity changes and is difficult to obtain high-precision SSP prediction results. In fact, the SSP estimation methods, that rely on a single data modality (sound field data or historical SSP data), are susceptible to data quality issues, such as high sound field measurement noise or low sound velocity sampling time resolution. To improve the robustness of the sound velocity estimation model, researchers have proposed some multimodal data fusion methods for estimating sound velocity that combining data from different sensors and sources to obtain more comprehensive features than a single data source [30, 21]. Yu et al. proposed a radial basis function (RBF) neural network for SSP estimation that mainly using historical temperature, salinity profile data and average SSP data. Nevertheless, the model is not sensitive to varieties of sound velocity, and the estimation results often approach the average SSP profile, making it difficult to accurately describe the spatiotemporal distribution changes of sound velocity. Xu et al. proposed a physics-inspired SOM model for SSP estimation, which introduced remote sensing SST data. However, SOM can only capture the influence of sea surface temperature on surface sound velocity distribution, and lacks the ability to capture large-scale feature correlations. Therefore, the accuracy of sound velocity estimation is difficult to meet practical application requirements.
To obtain real-time and accurate estimation of sound velocity distribution without underwater on-site data measurement, we fully consider the historical sound velocity distribution patterns of different spatial coordinates, as well as the impact of real-time SST changes on the dynamic characteristics of sound velocity distribution, and propose an interpretable SA-MDF-CNN model. In this model, the local correlation of features will be captured through CNN and global correlation of features will be captured through attention mechanism.
III SA-MDF-CNN Structure for SSP Estimation
To realize real-time and accurate estimation of sound velocity distribution, we propose an SSP estimation structure based on SA-MDF-CNN, which is shown in figure 1. In this paper, the ocean area of 5.5N-45.5N and 150.5E-170.5E are selected as the research scope. The remote sensing SST data, latitude and longitude coordinates, and the first three order feature vectors of historical SSPs decomposed by EOFs are first fused to construct the fusion data as the SA-MDF-CNN training data. Then, the multimodal fusion data of 8 grids around the task region is combined to construct the target SSP through the proposed SA-MDF-CNN model. In the following part, we will provide detailed introductions to data sources, data fusion structures, and the composition of neural network model, separately.
III-A Data sources
Remote sensing SST data are provided by the National Oceanic and Atmospheric Administration (NOAA) [31] with a spatial grid resolution of 0.25 and a temporal resolution of 1 day. The SSP data are obtained from the Chinese Observation and Research Station fo Global Ocean Argo System (Hangzhou) [32] with a spatial grid resolution of 1 and a temporal resolution of 1 month. Based on the objective analysis method of gradient-dependent correlation scale optimal interpolation, the 3D grid data of the subsurface layer (5-2000 meters) that covering 179.5W to 179.5E and 89.5S to 89.5N was constructed, and the observation profile was vertically interpolated to 57 standard layers with unequal intervals.
III-B Fusion data construction
Figure 2: The data fusion structure for a single coordinate.
Figure 3: SA-MDF-CNN model.
In the selected ocean region, we grid the entire area into N longitude-scales of 1, M latitude-scales of 1, and vertical depth of H layers. For any coordinate in the grid, the grid remote sensing SST data are expressed as $T^{s}_{n,m}$, and the monthly average SSP data of the historical period are chosen for data fusion, with the depth of the $h$th layer expressed as $D_{n,m}^{h},n\in N,m\in M$.
Within the divided $N*M$ grid ocean area, each sub-grid $\psi$ with a size of $3*3$ forms a set of training data, and each set of training data will be further divided into input data $X$ and output label data $Y$, with a sliding step of 1 for each sub grid, as shown in Figure 1. Specifically, for a sub-grid $\psi$, the fusion data formed by 8 surrounding coordinates will be the input data $X$, and the center SSP of each sub-grid $\psi$ will be taken as the output label $Y$. If the coordinate of the grid center is $L_{Y}=L_{n,m}$, then the coordinates of the input data are represented as:
$\mathcal{L_{X}}=[$ $L_{n-1,m-1},L_{n-1,m},L_{n-1,m+1},L_{n,m-1},L_{n,m+1},$ (1)
$L_{n+1,m-1},L_{n+1,m},L_{n+1,m+1}],\quad n\in N,\quad m\in M.$
For a single coordinate, the data fusion structure is shown in figure 2, which includes the real-time SST data, latitude and longitude coordinates, and the first three EOF feature vectors of the historical SSPs. EOF decomposition, also known as principal component analysis, is a statistical method used to analyze structural features in matrix data and extract major data feature quantities. Through EOF decomposition, the change information of the original sound velocity field can be easily condensed into the first few principal components and their corresponding spatial functions, so the main information of the sound velocity field can be expressed only by the first few eigenvectors.
To obtain the EOF feature vectors, corresponding 5-year (2015 to 2019) historical SSPs from the target ocean area are first selected and then linearly interpolated with a step size of 1 m. This standardized interpolation is used to more clearly describe the differences in sound velocity values at different depths. Suppose there are $J$ SSPs, and the $j$th SSP is expressed as a vector $\boldsymbol{S_{j}}=[s_{j,1},s_{j,2},...,s_{j,h}]^{T},h=1,2,...,H,H=1976$, where $s_{j,h}$ means the sound velocity value at the $h$th depth layer of the $j$th SSP. Since the SSP has been linearly interpolated, the depth of the $h$th depth layer is h m. Then, the matrix formed by all empirical SSPs will be:
$\mathcal{S_{H,J}}=\begin{bmatrix}s_{1,1}&s_{2,1}&\cdots&s_{J,1}\\ s_{1,2}&s_{2,2}&\cdots&s_{J,2}\\ \vdots&\vdots&\ddots&\vdots\\ s_{1,h}&s_{2,h}&\cdots&s_{J,h}\end{bmatrix}.$ (2)
Based on equation (2), the average SSP $\boldsymbol{S_{0}}=[s_{0,1},s_{0,2},...,s_{0,h}]^{T}$ can be calculated by averaging each row of equation (2). Next, a residual matrix can be constructed by subtracting $\boldsymbol{S_{0}}$ from each column of equation (2):
[…]
where $\mathcal{C_{H,H}}$ is a matrix with orders of $H\times H$. Performing EOF decomposition on $\mathcal{C_{H,H}}$ yields eigenvectors and eigenvector coefficients:
[…]
For the $j$th SSP in $\mathcal{S_{H,J}}$, it can be recovered by:
$\boldsymbol{S_{j}}=\boldsymbol{S_{0}}+\sum_{k=1}^{K}\alpha_{k}\boldsymbol{e_{k}},$ (6)
[…]
Finally, by fusing data from 8 coordinate points, input training data can be obtained. For location $L_{n-1,m-1}$, the fused data input $\mathcal{F}^{X}_{n-1,m-1}$ can be expressed as:
$\mathcal{F}^{X}_{n-1,m-1}=[\boldsymbol{T^{S}_{n-1,m-1}},\boldsymbol{L_{n-1,m-1}},\mathcal{E_{H,K}}],$ (8)
where $\boldsymbol{T^{S}_{n-1,m-1}}=[T^{s}_{n-1,m-1,1},T^{s}_{n-1,m-1,2},...,T^{s}_{n-1,m-1,H}]^{T}$ is actually a vector composed of H copies of $T^{s}_{n-1,m-1}$ at location $L_{n-1,m-1}$, and $\boldsymbol{L_{n-1,m-1}}$ is similar that $\boldsymbol{L_{n-1,m-1}}=[L_{n-1,m-1,1},L_{n-1,m-1,2},...,L_{n-1,m-1,H}]^{T}$. The label output data is $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$ at location $L_{n,m}$.
III-C SA-MDF-CNN Model
The network structure of SA-MDF-CNN is shown in Figure 3, mainly including attention module and convolutional network module. The first layer is the input layer, the size of which is 1976*6*8, including 6 types of physical quantities (remote sensing SST, latitude, longitude, and the first 3 order eigenvectors) at 8 coordinates with a total depth of 1976 meters. The second layer is a flattening layer, which flattens the three-dimensional fusion data into one-dimensional sequence data as input for the self-attention layer. The third layer comes the self-attention layer, which converts the fusion data into query, key and value vectors through three linear transforms. Then, the dot product of each query vector and each key vector is calculated to obtain the attention score matrix, which is further converted into attention weights through softmax function. Finally, the value vectors are summed using attention weights to obtain the final output features. These layers are designed to enhance the model’s ability of paying attention to the relationship between SSPs in different ocean areas in the input fusion data, so as to better capture the global dependence and key information between fusion data and the sound velocity distribution in the target ocean area.
The fourth and fifth layers are functional layers designed to reshape one-dimensional sequence data into two-dimensional data for subsequent feature extraction through convolution. The convolution layer is designed to further extract the features from the output data of the attention layer, which is better at capturing small-scale relationships of input features. The pooling layer is used to reduce the spatial dimension of the data, thereby reducing the number of parameters and reduce computational complexity, while retaining important feature information. The final two layers are the fully connected layer and the regression layer. The fully connected layer integrates the local features into the global features and outputs them to the regression layer. The RMSE is designed to be the loss function for calculating the gradient of the model parameters and updating weight parameters of the model through back propagation (BP) algorithm:
$Loss=\sqrt{\frac{\sum_{h=1}^{H}(\hat{s}^{Y}_{n,m,h}-s^{Y}_{n,m,h})^{2}}{H}},$ (9)
where $\hat{s}^{Y}_{n,m,h}$ is the estimated sound velocity value at depth $h$ of location $L_{n,m}$, and the estimated SSP can be expressed as $\hat{\mathcal{S}}^{Y}_{n,m}=[\hat{s}^{Y}_{n,m,1},\hat{s}^{Y}_{n,m,2},...,\hat{s}^{Y}_{n,m,H}]^{T}$. The SSP estimation algorithm based on SA-MDF-CNN is provided in Algorithm 1.
Algorithm 1 The SSP estimation algorithm based on SA-MDF-CNN.
1:Historical average SSP data $\mathcal{S}_{\mathcal{L_{X}}}^{mon}$, grid remote sensing SST $\mathcal{T}^{s}_{\mathcal{L_{X}}}$, position $\mathcal{L_{X}}$;
2:Initialize the parameters $\theta$, learning rate $\gamma$, and set the network parameters according to Table I;
3:for $k=1$ to $MaxEpoch$ do
4:     for $t=1$ to $maxBatchSize$ do
5:         Select a fusion data sample
6:    $\mathcal{F}^{X}_{n-1,m-1}=[\boldsymbol{T^{S}_{n-1,m-1}},\boldsymbol{L_{n-1,m-1}},\mathcal{E_{H,K}}]$;
7:         $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$;
8:         $Loss_{t}=\sqrt{\frac{\sum_{h=1}^{H}(\hat{s}^{Y}_{n,m,h}-s^{Y}_{n,m,h})^{2}}{H}}$;
9:         $\hat{\theta}\leftarrow\text{Adam}(\nabla_{\theta}\frac{1}{BatchSize}\sum Loss_{t},\theta,\gamma)$;
10:     end for
11:end for
12:Complete training and save SA-MDF-CNN model;
13:Estimate and output target SSP $\mathcal{S}_{n,m}^{es}$ at intermediate coordinates $L^{test}_{n,m}$.
III-D Multi-head self-attention module
Figure 4: The principle of the multi-head self-attention mechanism.
The self-attention mechanism is a special type of attention mechanism used to process sequential data such as text, images, or time series. It allows the model to process the sequence in a way that considers the relationship between each element in the sequence and all other elements. The self-attention mechanism calculates weights by evaluating the similarities between different elements in a sequence to determine how much an element is related to other elements. Specifically, a kind of similarity score is defined to describe the similarity between each element in the sequence and other elements. Then, based on these similarity scores, the weight of each element can be calculated, and a weighted representation can be further obtained by multiplying the weight with the corresponding element. In order to better capture the associations between different types of SSP, multiple attention heads are used in this paper as shown in figure 4.
The diagram presented in Figure 4 illustrates the intricate workings of the multi-head self-attention mechanism, which is a fundamental building block in the architecture of transformer models. Three pivotal components are involved in the self-attention mechanism: the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ vectors. These vectors are derived from the input data through linear transformations, as depicted in the upper right section of the diagram where the input is projected into the query, key, and value spaces using learnable weight matrices.
The query vector serves to assess the relevance or correlation with the key vectors. This correlation is quantified and utilized to compute a weighted sum, which amalgamates the contributions of the value vectors. The key vectors are instrumental in determining the attention distribution function, denoted as $a_{i}$ in the diagram, which signifies the importance or weight of each element within the input sequence. In contrast, the value vectors encapsulate the specific features or numeric values associated with each element, playing a crucial role in the aggregation of information.
The mechanism operates by first computing the dot product between the query and each key vector. To prevent the dot products from growing too large and causing numerical instability, especially when the dimensionality of the keys $d_{k}$ is high, each key is normalized by dividing it by $\sqrt{d_{k}}$. This normalization step is shown in the diagram as part of the attention calculation process. Following this, the softmax function is applied to the normalized dot products to generate a set of attention weights. These weights are then used to compute a weighted sum of the value vectors, as represented by the equation:
$\text{Attention}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}$ (10)
This operation is performed for each head in the multi-head self-attention mechanism, as shown in the lower section of the diagram. Each head attends to different parts of the input sequence, allowing the model to capture a diverse range of dependencies and nuances within the data. The outputs from all heads are then concatenated and subjected to a linear transformation, resulting in the final output of the multi-head self-attention mechanism.
Given an input vector $\mathbf{x}\in\mathbb{R}^{dim}$, represented as $\mathcal{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$, where $dim$ denotes the number of feature dimensions and $M$ represents the number of sequence samples. Upon receiving a query vector $\mathbf{q}$, the self-attention function is defined as:
$\mathrm{att}(\mathcal{X},\mathbf{q})=\sum_{i=1}^{M}a_{i}\mathbf{x}_{i}=\sum_{i=1}^{M}\frac{\exp(\mathrm{s}(\mathbf{k}_{i},\mathbf{q}))}{\sum_{j=1}^{M}\exp(\mathrm{s}(\mathbf{k}_{j},\mathbf{q}))}\mathbf{v}_{i},$ (11)
where $a_{i}$ is the $i$th attention weight, computed by the softmax function over the scaled dot product of the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$. The function $\mathrm{s}(\mathbf{k}_{i},\mathbf{q})$ represents the similarity score between the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$, which is typically calculated as the dot product $\mathbf{k}_{i}^{\top}\mathbf{q}$ divided by the square root of the key vector’s dimension $\sqrt{d_{k}}$. The value vector $\mathbf{v}_{i}$ corresponds to the $i$th sample in the input sequence, and the output of the self-attention function is a weighted sum of these value vectors, where the weights are determined by the attention mechanism.
In a standard self-attention module, the multi-head attention mechanism employs multiple query vectors $\mathcal{Q}$ to select multiple sets of information from the input data in parallel. Specifically, each head independently computes the attention weights between the query vectors $\mathcal{Q}$, key vectors $\mathcal{K}$, and value vectors $\mathcal{V}$ to extract different subsets of information. The output of the multi-head self-attention module is calculated by concatenating the outputs of all heads and then applying a linear transformation. The mathematical expression is:
$\mathrm{MultiHead}(\mathcal{Q},\mathcal{K},\mathcal{V})$ $=\mathrm{Concat}(\mathrm{head}_{1},\ldots,\mathrm{head}_{\mathrm{h}})\mathcal{W}^{O},$ (12)
where $\mathrm{head_{i}}=\text{Attention}(\mathcal{QW}_{i}^{Q},\mathcal{KW}_{i}^{K},\mathcal{VW}_{i}^{V})$, $\mathcal{W}_{i}^{Q}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{K}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{V}\in\mathbb{R}^{d_{v}}$, and $\mathcal{W}^{o}\in\mathbb{R}^{hd_{v}}$ is a projection of the parameter matrix.
IV Results and discussions
To test the effectiveness of the proposed SA-MDF-CNN model, the monthly average fusion data from July 2015 to 2019 were selected as input training data and the monthly average SSP from 2020 were set to be the output label data. Experimental results were compared with other widely used methods for constructing spatial sound velocity distribution, including convolutional neural network (CNN), spatial interpolation (SITP) and mean values (MEAN), respectively. The model parameter settings are given in Table I.
TABLE I: Parameter settings of SA-MDF-CNN
Parameter Value
GPU RTX 3090
input size [1976,6,8]
self-attention [8,32]
filter size [2,2]
number of Channels 256
number of filters 256
convolution stride size [1,1]
pool size [2,2]
pooling stride 2
full connected output size 1976
minimum batch size 16
maximum epochs 100
learning rate $10^{-3}$
drop factor of learning rate 0.1
drop period of learning rate 20
IV-A Accuracy Performance
Figure 5: Remote sensing SST data.
To test the accuracy performance of proposed model in estimating SSPs, real-time remote sensing SST data from different locations and depths were fused with EOF feature vectors to construct the sound velocity field. The remote sensing SST data are shown in figure 5, and a visual display of the fusion data is given in figure 6.
Figure 6: Fusion data of remote sensing SST data, latitude and longitude coordinates, and EOF feature vectors.
Figure 7: Comparison of real-time estimated SSP results of different algorithms with depth of 1976 m, where (a) is the SA-MDF-CNN algorithm, (b) is the CNN algorithm, (c) is the spatial interpolation algorithm, and (d) is the mean value method.
TABLE II: Comparison of RMSE at the same depth:1976 meters.
Number Location RMSE of different algorithms (m/s)
SA-MDF-CNN CNN SITP Mean method
1 7.5°N 156.5°E 0.1125 0.1317 0.2456 0.2591
2 7.5°N 151.5°E 0.2599 0.3084 0.3182 0.3377
3 6.5°N 157.5°E 0.1293 0.1402 0.1730 0.1812
4 8.5°N 163.5°E 0.1152 0.1219 0.2775 0.2903
5 11.5°N 165.5°E 0.2090 0.2149 0.2709 0.2842
6 24.5°N 162.5°E 0.0787 0.1115 0.1232 0.1281
7 26.5°N 163.5°E 0.1181 0.1378 0.1293 0.1367
8 27.5°N 153.5°E 0.1039 0.1239 0.1532 0.1621
9 7.5°N 160.5°E 0.1945 0.2050 0.4025 0.4229
Average 0.1468 0.1661 0.2326 0.2447
The real-time SSP estimation results of SA-MDF-CNN algorithm with depths ranging from 0 to 1976 meters within $7^{\circ}N$ to $28^{\circ}N$, $150^{\circ}E$ to $165^{\circ}E$ are compared with those of CNN, SITP and mean values method as shown in figure 7. From the intuitive representation of the sound velocity heat-map, it can be seen that the sound velocity distribution estimated by SA-MDF-CNN and CNN has a significant improvement in accuracy compared to SITP and Mean value methods, and the estimated values of the proposed model are more similar to the sound velocity distribution of real samples. A comprehensive evaluation of the accuracy performance of various algorithms across different locations at depths ranging from 0 to 1976 meters is presented in Table II. Intuitive observation shows that the average RMSE results of SA-MDF-CNN, CNN, SITP, and Mean value methods are 0.1468, 0.1661, 0.2326, and 0.2447, respectively. Notably, the RMSE of SA-MDF-CNN for real-time SSP estimation is approximately 12% lower than that of CNN and approximately 30-40% lower than that of traditional SITP and mean value methods. Figure 8 gives a comparison of sound velocity estimation curves at $7.5^{\circ}N,156.5^{\circ}E$. In figure 8 (a), the curve of SA-MDF-CNN is the closest to the real SSP, and in figure 8 (b), the error mainly manifests in the shallow ocean area within 500 meters because the sound velocity in shallow water is more severely and irregularly affected by temperature. However, among all methods the error disturbance of SA-MDF-CNN is the smallest.
Figure 8: A comparison example of real-time estimated SSP results of different algorithms with depths ranging from 0 to 1976 meters, where (a) is the comparison of estimated results and real SSP, and (b) is the comparison of error distribution between the estimated results of different algorithms and the real SSP.
Figure 9: Comparison of estimation results of different algorithms at different shallow ocean depths.
TABLE III: Comparison of RMSE at different depth.
Depth Location RMSE of different algorithms (m/s)
SA-MDF-CNN CNN SITP Mean method
200m 6.5°N 159.5°E 0.6075 0.6911 1.085 1.143
7.5°N 156.5°E 0.2911 0.3468 0.6688 0.7035
7.5°N 161.5°E 0.3859 0.4092 1.019 1.085
9.5°N 164.5°E 0.1383 0.1775 0.3831 0.4018
13.5°N 169.5°E 0.3115 0.3556 0.5565 0.5862
Average 0.3469 0.3960 0.7425 0.7839
300m 11.5°N 165.5°E 0.4957 0.5155 0.6903 0.7240
12.5°N 152.5°E 0.1901 0.3242 0.3392 0.3542
13.5°N 154.5°E 0.2889 0.3028 0.4363 0.4594
14.5°N 158.5°E 0.1360 0.1411 0.3860 0.4088
14.5°N 166.5°E 0.2943 0.3446 0.5869 0.6156
Average 0.2810 0.3256 0.4877 0.5124
500m 7.5°N 151.5°E 0.2293 0.2539 0.4839 0.5101
6.5°N 157.5°E 0.2299 0.2613 0.3371 0.353
8.5°N 164.5°E 0.2943 0.3257 0.6033 0.6368
11.5°N 156.5°E 0.1571 0.1625 0.3331 0.3481
17.5°N 163.5°E 0.1133 0.1841 0.2544 0.2721
Average 0.2048 0.2375 0.4024 0.4240
Figure 10: Comparison of error distributions of real-time estimated SSPs for different algorithms at different shallow sea depths.
To further test the accuracy and performance of the model in estimating the sound velocity distribution in shallow water areas, the real-time SSP estimation results of SA-MDF-CNN algorithm at 200 m, 300 m and 500 m depths are compared with those of CNN, SITP and mean values method as shown in figure 9. As can be seen from the enlarged part of each sub-figure, the estimated SSP curves by SA-MDF-CNN at 200 m, 300 m and 500 m depths are significantly closer to the real SSP prediction results than those of CNN, SITP and mean method at the same ocean area. Specifically, an in-depth investigation into the performance of different algorithms in shallow ocean areas at varying depths and locations was conducted, with the real-time SSP estimation RMSE results compared and displayed in Table LABEL:table3. The largest RMSE is observed at a depth of 200 meters, which can be attributed to the increased interference from real-time sea conditions in the shallow sea environment, such as wind speed, wave height, tsunami, and other factors. Concurrently, the experimental results indicate that the estimation accuracy of the proposed SA-MDF-CNN model at 200 meters depth is enhanced by approximately 13% compared to CNN and by approximately 53% compared to SITP and mean value method. At an ocean depth of 300 meters, the proposed algorithm demonstrates an improvement of about 14% over CNN and about 42% over SITP and mean value method. At 500 meters depth, the proposed model shows an improvement of about 14% compared to CNN and about 49% compared to SITP and mean value method. These experimental results collectively illustrate that the proposed model maintains significant precision advantages even in shallow sea environments of varying depths, thereby underscoring its applicability and effectiveness. The results indicate that SA-MDF-CNN is also applicable even in shallow ocean environment with different depths, and still has smaller RMSE compared with other algorithms, which means that the model still has good feature capture ability in shallow ocean scenes with significant changes in sound velocity. The average absolute error distribution comparison of real-time estimated SSP of different algorithms tested in the set of 200m and 500m depths is given in Figure 10. The error fluctuation of SA-MDF-CNN is the most stable, which indicates that SA-MDF-CNN still has better robust performance under shallow sea conditions with different depths.
Figure 11: Interpretability analysis of the SA-MDF-CNN model, where (a) is the first three order EOF the historical SSP at different positions (b) is the attention weight with Epoch=10, (c) is the attention weight with Epoch=50, and (d) is the attention weight at Epoch=10.
IV-B Interpretability analysis
To investigate why the proposed SA-MDF-CNN model has better performance compared to CNN, we conducted an experiment and visually represented the attention weights of the model as shown in figure 11. Figure 11 (a) illustrates the first three order EOF feature vectors of the historical SSPs at various latitude and longitude coordinates. Figure 11 (b), (c), and (d) correspond to the attention weights when the training epochs are 10, 50, and 100, respectively. Through observing the evolution trend of attention focus with increasing training times, we found that the SA-MDF-CNN model increasingly focuses on the sound velocity distribution in shallow waters within 1000 meters. This trend aligns with the first three orders of EOF feature distribution shown in panel (a), although the EOF itself emphasizes the SSP distribution between 0-500 meters. The attention depth of the proposed model spans 0-1500 meters, which is attributed to the variation of SSP distribution under different latitude and longitude coordinates. Therefore, in order to enhance the model’s ability to focus on the global SSP distribution, we added a multi-head self-attention layer after input data fusion. This mechanism further confirms the effectiveness of the proposed model. Experiments based on the aforementioned public dataset demonstrate that the proposed SA-MDF-CNN model outperforms other algorithms in estimating sound velocity in both shallow and deep waters.
IV-C Efficiency and stability
TABLE IV: Comparison of Neural Networks
Network Number of Parameter Train time
CNN 61.9M 72s
SA-MDF-CNN 149M 40s
Although the accuracy performance in SSP estimation of SA-MDF-CNN is better than traditional methods, the disadvantage is that SA-MDF-CNN increases the number of parameters and computational overhead. Table IV shows the comparison of parameters and running time between the proposed model and CNN. It can be seen that the proposed model in this paper needs more parameters and training time than CNN, which is a necessary overhead for performance improvement.
Figure 12: Comparison of RMSE convergence of different algorithms.
To test the stability of the model, the convergence of SA-MDF-CNN and CNN with different numbers of iteration is given in figure 12. Both SA-MDF-CNN and CNN can converge in less than 200 training iterations. The convergence of SA-MDF-CNN is smoother than that of CNN, which may be due to the multi-head self-attention mechanism focusing more on global features and therefore not causing significant parameter updates due to drastic changes in a small range of the data.
V Ocean experiments
Figure 13: Data sampling locations of ocean experiments.
Figure 14: Comparison of real-time SSP estimation results of different algorithms on ocean experimental data, where (a) shows the SSP curves, and (b) shows the error distributions.
In order to verify the practicality of the SA-MDF-CNN model, a comprehensive deep-ocean experiment was conducted in the South China Sea in April 2023. As illustrated in figure 13 (a), various instruments such as CTD and XCTD were employed at diverse marine locations, yielding a total of 14 distinct samples. The depth of sampling reached 3500 meters, with intervals set at every 100 meters. It is evident that the selected location for this experiment does not conform to the rigid 3x3 grid pattern, thus allowing for a more rigorous evaluation of the proposed model. The fusion data at the blue coordinates in figure 13 (b) were selected as the input of the model, and the real-time SSP under the red pentagram was used as the prediction output of the model.
Figure 15: Comparison of real-time estimation SSP results of different algorithms in shallow sea environment with different depths in ocean experiment, where (a) and (c) are SSP curves of 300 meters depth, (b) and (d) are error distributions of 500 meters depth.
Figure 16: RMSE comparison of real-time estimated SSPs by different algorithms of ocean experiments.
To evaluate the effectiveness of the model in estimating the distribution of sound velocity at full-ocean depth, the estimated SSP of SA-MDF-CNN is compared with other algorithms in figure 14. It is evident that the proposed model exhibits lower RMSE and more stable error disturbance compared to traditional algorithms and CNN. To further test the sound velocity estimation performance in shallow waters with significant changes in sound velocity, estimation of sound velocity at 300, 500 depth meters by SA-MDF-CNN are compared with other methods in figure 15. The results indicate that the proposed algorithm has smaller error disturbances and maintains significant accuracy advantages compared to other algorithms in shallow water environments. Figure 16 gives a more detailed comparison of the RMSE results of different algorithms, clearly showing that the proposed model has significant advantages at different depths.
VI Conclusion
To construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, a SA-MDF-CNN model driven by multimodal data fusion is proposed in this paper, which fuses historical SSP features, real-time remote sensing SST, latitude and longitude coordinates. The fusion data is constructed by gridding, and the real-time SSP is retrieved based on the real-time remote sensing SST. The effectiveness of our proposed model was tested in the case of non-standardized grid through ocean experiments. The experimental results show that our proposed model is not only suitable for real-time SSP estimation under non-strictly standardized grid conditions, but also suitable for shallow sea conditions with different depths.
Acknowledgments
The authors acknowledge the historical SSP data support from the China Argo Real-time Data Center, (https://www.argo.org.cn/, latest access:December 20, 2024). The remote sensing SST data can be acquired at the National Oceanic and Atmospheric Administration (https://www.commerce.gov/, latest access: December 25, 2024).

### v3

@@
-An Attention-Assisted AI Model for Real-Time Underwater Sound Speed Estimation Leveraging Remote Sensing Sea Surface Temperature Data
+An Attention-Assisted Multi-Modal Data Fusion Model for Real-Time Estimation of Underwater Sound Velocity
 Abstract
@@
 The ultimate objective is to facilitate a rapid and precise estimation of sound velocity distribution within a specified task area.
-The comparative analysis demonstrates that the proposed approach achieves superior performance in terms of both accuracy and stability, exhibiting reduced error rates and enhanced resistance to disturbances when benchmarked against existing advanced techniques.
+The comparative analysis demonstrates that the proposed approach achieves superior performance in terms of both accuracy and stability against existing advanced techniques.
 Index Terms:
@@
 Underwater sound velocity construction plays an indispensable role in many applications such as underwater positioning, navigation, timing, communication (PNTC) and target recognition due to its decisive factor in the signal propagation mode [1, 2, 3].
-Acoustic propagation speed in marine settings exhibits spatial variability due to the combined effects of temperature, salinity, and pressure [4, 5, 6, 7].
-Generally, the sound velocity exhibits a stratified variation with depth, so that sound speed profiles (SSPs) are commonly employed to describe the distribution of sound velocity.
+Acoustic propagation speed in marine environment exhibits spatial variability due to the combined effects of temperature, salinity, and pressure [4, 5, 6, 7].
+Generally, the sound velocity exhibits a stratified variation with depth within the kilometer level range, so that sound speed profiles (SSPs) are commonly employed to describe the distribution of sound velocity.
 The methodology for acquiring SSPs has been a popular research topic for a long period.
-In traditional ways, SSPs can be directly measured using various instruments, including the conductivity–temperature–depth profiler (CTD), the expendable CTD (XCTD), and the sound velocity profiler (SVP).
+In traditional ways, SSPs can be directly measured by various instruments, including the conductivity–temperature–depth profiler (CTD), the expendable CTD (XCTD), and the sound velocity profiler (SVP).
 Alternatively, SSPs can be derived by inversion based on acoustic field data, which encompasses parameters such as signal transmission time or received signal strength [8, 9, 10, 11, 12].
 In general, instrument-based measurements can provide precise SSPs with high–depth resolution, but the measurement process is notably time-intensive.
-For instance, when the instrument is deployed at a regular speed of 50 meters per minute, it takes at least 80 minutes to measure the sound velocity value within a depth range of 2000 meters to complete the deployment and retrieval of the equipment [11].
+For instance, when the instrument is deployed at a regular speed of 50 meters per minute, it takes at least 80 minutes to measure the sound velocity value within a depth range of 2000 meters, including the deployment and retrieval of the equipment [11].
 As a result, it is difficult to provide real-time sound velocity distributions for underwater communication and positioning through instrument-based measurement.
-Considering that the distribution of sound velocity can affect the distribution of sound field, researchers have proposed a series of methods to invert the sound velocity distribution by using on-site measured sound field data such as signal propagation time in order to accelerate the process of obtaining sound velocity.
+Since the distribution of the sound speed affects the distribution of the acoustic field, scholars in related areas have come up with a range of approaches to invert the sound speed distribution.
+They make use of the on-site measured data of the sound field such as the signal propagation time, aiming at accelerating the process of obtaining the sound velocity distribution.
 There are three main frameworks for inversion methods, including matching field processing (MFP) [13], compressive sensing (CS) [14, 15], and deep learning (DL) [4, 16].
-The proposed inversion technique enables fast acquisition of sound speed profiles, offering substantial temporal advantages over conventional conductivity-temperature-depth (CTD) sensors and sound velocity profilers (SVPs) in marine data collection.
+The proposed inversion technique enables fast acquisition of sound speed profiles, offering substantial temporal advantages over conventional CTD sensors and SVPs in marine data collection.
 However, because of the interference of noise in sound field measurement, there is a loss in the accuracy of these sound velocity inversion methods.
@@
 Over the past few years, the fast development of underwater automated sensing equipment has provided increasingly rich vertical observation reference data for exploring the distribution of underwater sound velocity.
-To achieve fast estimation of SSPs and eliminate the need for underwater on-site operations, various methods have emerged that use empirical sound velocity distribution data or remote sensing data.
-[17, 18, 19] utilize empirical SSPs data to capture the changes in sound velocity from a time series perspective, enabling real-time estimation and prediction of sound velocity distribution in designated ocean areas.
+To achieve fast estimation of SSPs and eliminate the need for underwater on-site operations, various methods have emerged that use historical sound velocity distribution data or remote sensing data.
+[17, 18, 19] utilize historical SSPs data to capture the changes in sound velocity from a time series perspective, enabling real-time estimation and prediction of sound velocity distribution in designated ocean areas.
 These SSP prediction methods have a single type of input data and low complexity of model construction.
 However, because of the low time resolution of historical samples, its accuracy performance is not good enough in estimating the sound velocity within a small-scale time range.
-In fact, the variation of sound velocity distribution in a small-scale spatial area is mainly reflected in the shallow water part, as it is more significantly affected by temperature changes.
+In fact, the temporal variation of sound velocity distribution in small-scale spatial regions is mainly reflected in the shallow water part, as it is more significantly affected by temperature changes.
 Marine remote sensing technology provides real-time and reliable high spatiotemporal resolution sea surface temperature (SST) data [20], which can offer initial sea surface condition constraints for estimating underwater sound velocity distribution.
-[21] proposed a self-organizing map (SOM) neural network model that combines empirical sound velocity data and remote sensing SST data, but the model mainly focuses on the local characteristics of sound velocity distribution, which not only fails to capture the influence of SST data on deep ocean sound velocity, but also fails to capture the spatial correlation of sound velocity distribution.
-As a result, the accuracy performance has not been significantly improved compared to prediction methods.
+Xu et al [21] introduced an innovative self-organizing map (SOM) architecture that integrates empirical acoustic propagation measurements with satellite-derived SST observations to establish robust correlations between marine acoustic properties and surface thermal patterns, but the model mainly focuses on the local characteristics of sound velocity distribution, which not only fails to capture the influence of SST data on deep ocean sound velocity, but also fails to capture the spatial correlation of sound velocity distribution.
+As a result, the accuracy performance is difficult to meet the application requirements of high-precision PNTC systems.
 In order to fast and accurately estimate the SSP of a given task area without on-site underwater data measurement, we propose an interpretable self-attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model in this paper to deeply capture the intrinsic relationships among empirical SSP data, remote sensing SST data, and spatial coordinates.
-The core idea is to capture the local feature correlation of regional sound velocity distribution through CNN, and enhance the learning ability of global feature correlation of multimodal data through attention mechanism.
-To evaluate the effectiveness of the proposed SA-MDF-CNN model, we conducted experiments using Argo data from the Pacific Ocean and measured data from the South China Sea in 2023.
-Results show that SA-MDF-CNN can achieve lower root mean square error (RMSE) than other spatial SSP construction methods.
-The contribution of this paper can be summarized as follows:
-  • To achieve real-time estimation of SSP without on-site underwater data measurement, we propose an attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model.
-We take into account the mutual influence among empirical sound velocity distribution, SST, and spatial position, and the attention module is designed to enhance the global feature learning ability of the model.
+The core idea focuses on extracting spatial patterns of regional sound velocity interdependencies via convolutional neural networks, while the application of attention mechanisms enhances the model’s capacity to learn complex interactions within multimodal data.
+In an attempt to comprehensively appraise the overall effectiveness of the purportedly proposed SA-MDF-CNN model, we carried out a series of carefully designed experiments by utilizing the Argo data originating from the expanse of the Pacific Ocean and the precisely measured data obtained from the South China Sea in the year 2023.
+The detailed results clearly indicate that the SA-MDF-CNN model has the distinct ability to accomplish a notably lower root mean square error (RMSE) value compared to various other existing spatial SSP construction methods.
+The significant contributions presented in this research paper can be summarized in the following ways:
+  • To achieve real-time estimation of SSP without on-site underwater data measurement, we propose an interpretable SA-MDF-CNN model.
+We take into account the mutual influence among historical sound velocity distribution, SST, and spatial position, and the attention module is developed to optimize the model’s comprehensive feature extraction capabilities across spatial domains.
   • To evaluate the accuracy performance of the model, we conducted experiments using Argo data from the Pacific region and further validated it through sea trials in the South China Sea in 2023.
 The RMSE of both experimental results is superior to other spatial SSP constructing methods.
-  • To enhance the interpretability of the model, the weights of the model parameters are visualized.
-As the number of training iterations increases, the model parameter weights gradually focus on the shallow water part.
-This is reasonable because through the empirical orthogonal function (EOF) decomposition of historical SSPs, it can be seen that the differences in sound velocity distribution are more prominent in shallow water.
-Therefore, the weight visualization results indicate that the model can effectively capture the characteristics of sound velocity distribution.
-The structure of this article is arranged as follows.
-Section 2 provides a detailed introduction to relevant works.
-The overall architecture and functional module description of the SA-MDF-CNN model are provided in Section 3.
-Section 4 gives the experimental findings, and the final section, Section 5, presents the concluding remarks.
+  • To boost the model’s interpretability, a visualization of the model parameter weights is carried out.
+With the escalation of training iterations, the weights of model parameters progressively concentrate on the shallow water region.
+This phenomenon is justifiable as, upon performing empirical orthogonal function (EOF) decomposition on historical SSPs, it becomes evident that disparities in sound velocity distribution are more pronounced in shallow water.
+Consequently, the outcomes of weight visualization suggest that the model is capable of efficiently capturing the traits of sound velocity distribution.
+The organization of this paper is designed as follows.
+In Section 2, a brief elaboration on related works is offered.
+Section 3 puts forward the general framework and the descriptive details of the functional modules of the SA-MDF-CNN.
+The outcomes of the experiment are reported in Section 4.
+Concluding statements are presented in Section 5.
 II Related works
-The distribution of underwater sound velocity plays an essential role in the energy distribution and propagation trajectory of acoustic signals.
-Therefore, real-time and accurate estimation of sound velocity distribution has significant value for applications based on communication and positioning technologies such as underwater PNTC systems, and target recognition systems.
+The spatial structure of underwater sound velocity field plays a decisive role in the energy distribution and propagation trajectory of acoustic signals.
+Therefore, real-time and accurate estimation of sound velocity distribution is of great meaningful for applications based on communication and positioning technologies such as underwater PNTC systems, and target recognition systems.
 The traditional way of obtaining sound velocity distribution is usually through direct measurement using shipborne CTD or SVP equipment, which has the advantage of high accuracy [22, 23].
 However, it comes with high economic costs and measurement time expenses.
-With the development of sensing and network technology, many kinds of underwater environmental observation systems have been established to enhance human understanding of the ocean.
-In 1979, MUNK and Wunsch first put forward the concept of ocean acoustic tomography that using acoustic field observation information to invert the sound velocity distribution [24, 25].
+The development of sensor network technologies has led to the implementation of multiple underwater environmental monitoring frameworks, advancing scientific comprehension of oceanic processes.
+The methodological framework of underwater acoustic inversion was initially formalized in 1979 by Munk et al.[24, 25], introducing an innovative paradigm that leverages hydroacoustic propagation measurements for reconstructing SSPs.
 Since then, there have been three mainstream frameworks for sound velocity inversion, namely MFP [13], CS [14, 15], and DL [4, 16].
@@
 Then, different feature combinations are generated to form candidate SSPs.
-Next, the sound field distribution is simulated based on ray theory.
-Finally, the simulated sound field is matched with the measured sound field to determine the optimal estimation of SSP.
-To accelerate the search speed of the optimal candidate SSP, Tolstoy introduced simulated annealing algorithm in the MFP, which improved the algorithm execution efficiency but resulted in suboptimal solutions [13].
-Afterwards, other heuristic algorithms were combined with MFP frameworks for SSP inversion, but they also obtained suboptimal solutions [26, 27].
+Subsequently, acoustic field distributions are computationally modeled employing an acoustic ray tracing theory.
+The concluding phase involves cross-validation between the modeled sound field distributions and experimentally measurements to establish the optimal inversion parameters for the SSP.
+To accelerate the search process of the optimal candidate SSP, Tolstoy introduced simulated annealing algorithm in the MFP, which improved the algorithm execution efficiency but only suboptimal solutions were obtained [13].
+Afterwards, other heuristic algorithms were combined with MFP frameworks for SSP inversion [26, 27], but they have the same problem as [13].
 To further improve the efficiency of inversion algorithm execution, Bianco [15] and Choo [14] proposed SSP inversion framework based on CS, respectively.
-In the CS framework, the mapping relationship from the sound field distribution to the sound velocity distribution is established through a matrix, which eliminates the search process for matching terms.
+Within the context of the CS framework, the correlation between the sound field distribution and the sound velocity distribution is formulated via a matrix based approach, which eliminates the search process for matching terms.
 However, the matrix relationship introduces linear approximation, thus sacrificing inversion accuracy.
-In recent years, deep learning theory has developed rapidly and achieved significant results in constructing unknown nonlinear relationships between different data [28, 29].
+The theoretical frameworks of deep neural networks (DNNs) have experienced accelerated progression in contemporary research, particularly in establishing intricate nonlinear dependencies across heterogeneous data domains [28, 29].
 In addition, long-term underwater environmental observations have accumulated a large amount of data for marine hydrological research, laying a data foundation for the application of deep learning underwater.
-To address the drawbacks of computational efficiency and accuracy loss in MFP and CS frameworks, we proposed a auto-encoding feature mapping neural network model for SSP inversion in our early work [4], in which the auto-encoder was created to extract deep robust features so as to reduce the impact of acoustic field measurement errors on the accuracy of sound velocity inversion.
+To address the drawbacks of computational efficiency and accuracy loss in MFP and CS frameworks, an auto-encoding neural network model for feature mapping was introduced for SSP inversion in our prior research [4], in which the auto-encoder was created to extract deep robust features so as to reduce the impact of acoustic field measurement errors on the accuracy of sound velocity inversion.
 Considering the insufficient accumulation of historical sound velocity data in some ocean areas, the DL model is prone to over-fitting and reduces accuracy performance.
-Therefore, we further propose a few-shot meta learning framework to accelerate the convergence speed of the model [16].
-However, the above-mentioned methods based on MFP, CS, and DL all rely on real-time measured acoustic field data, which puts high demands on the deployment of underwater observation equipment.
+Therefore, we proposed a meta-learning framework tailored for few-shot scenarios to enhance the rate at which the model converges [16].
+However,the methodologies aforementioned, which are predicated on MFP, CS, and DL, all rely on real-time measured acoustic field data, which imposes strict requirements on the arrangement of underwater observation equipment.
 Therefore, these methods not only face high equipment economic costs, but also have limited application scope for areas without sonar measurement systems.
 Figure 1: SSP Estimation Structure based on SA-MDF-CNN.
-Nowadays, how to eliminate the need for underwater on-site data measurement has become a research hotspot in the field of SSP inversion.
-Based on this requirement, Liu et al.
-and Lu et al.
-both proposed a long-short term memory (LSTM) neural network model for SSP prediction [17, 18], which only requires empirical SSP data.
-However, due to the low temporal resolution of prior data in most marine areas, this prediction method can only describe the overall trend of sound velocity changes and is difficult to obtain high-precision SSP prediction results.
+Nowadays, removing the necessity for on-site underwater data measurement has emerged as a prominent research focus in the domain of SSP inversion.
+In accordance with this demand, both Liu et al.
+[17] and Lu et al.
+[18] put forward a Long-Short Term Memory (LSTM) neural network model for SSP forecasting, which solely requires historical SSP data.
+Nevertheless, owing to the insufficient temporal resolution of prior data in the majority of marine regions, this prediction method can only describe the overall trend of sound velocity changes and is difficult to obtain high-precision SSP prediction results.
 In fact, the SSP estimation methods, that rely on a single data modality (sound field data or historical SSP data), are susceptible to data quality issues, such as high sound field measurement noise or low sound velocity sampling time resolution.
-To improve the robustness of the sound velocity estimation model, researchers have proposed some multimodal data fusion methods for estimating sound velocity that combining data from different sensors and sources to obtain more comprehensive features than a single data source [30, 21].
+To improve the robustness of the sound velocity estimation model, scholars have proposed some multimodal data fusion methods for estimating sound velocity that combining data from different sensors and sources to obtain more comprehensive features than a single data source [30, 21].
 Yu et al.
-proposed a radial basis function (RBF) neural network for SSP estimation that mainly using historical temperature, salinity profile data and average SSP data.
-Nevertheless, the model is not sensitive to varieties of sound velocity, and the estimation results often approach the average SSP profile, making it difficult to accurately describe the spatiotemporal distribution changes of sound velocity.
+[30] proposed a radial basis function (RBF) neural network for SSP estimation that mainly using historical temperature, salinity profile data and average SSP data.
+Nevertheless, the model is not sensitive to varieties of sound velocity, and the estimation results often approach the average SSP profile, resulting in challenges in accurately depicting the spatiotemporal variations of sound velocity.
 Xu et al.
-proposed a physics-inspired SOM model for SSP estimation, which introduced remote sensing SST data.
-However, SOM can only capture the influence of sea surface temperature on surface sound velocity distribution, and lacks the ability to capture large-scale feature correlations.
+[21] proposed a physics-inspired SOM model for SSP estimation, which introduced remote sensing SST data.
+However, SOM is confined to capturing the impact of SST on the distribution of sound velocity at the surface, and lacks the ability to capture large-scale feature correlations.
 Therefore, the accuracy of sound velocity estimation is difficult to meet practical application requirements.
@@
 To realize real-time and accurate estimation of sound velocity distribution, we propose an SSP estimation structure based on SA-MDF-CNN, which is shown in figure 1.
-In this paper, the ocean area of 5.5N-45.5N and 150.5E-170.5E are selected as the research scope.
-The remote sensing SST data, latitude and longitude coordinates, and the first three order feature vectors of historical SSPs decomposed by EOFs are first fused to construct the fusion data as the SA-MDF-CNN training data.
-Then, the multimodal fusion data of 8 grids around the task region is combined to construct the target SSP through the proposed SA-MDF-CNN model.
+In this paper, the ocean area of 5.5N-45.5N and 150.5E-170.5E are selected as the research area.
+The remote sensing SST data, latitude and longitude coordinates, and principal component eigenvectors derived from EOF decomposition of historical SSPs are first fused to construct the fusion data as the SA-MDF-CNN training data.
+Then, the multimodal fusion data of 8 grids around the task region are combined through the proposed SA-MDF-CNN model to construct the target SSP.
 In the following part, we will provide detailed introductions to data sources, data fusion structures, and the composition of neural network model, separately.
@@
 Remote sensing SST data are provided by the National Oceanic and Atmospheric Administration (NOAA) [31] with a spatial grid resolution of 0.25 and a temporal resolution of 1 day.
+The SST data are arithmetic averaged on a monthly basis to achieve the consistency of the fused data in time sampling.
 The SSP data are obtained from the Chinese Observation and Research Station fo Global Ocean Argo System (Hangzhou) [32] with a spatial grid resolution of 1 and a temporal resolution of 1 month.
@@
 $L_{n+1,m-1},L_{n+1,m},L_{n+1,m+1}],\quad n\in N,\quad m\in M.$
-For a single coordinate, the data fusion structure is shown in figure 2, which includes the real-time SST data, latitude and longitude coordinates, and the first three EOF feature vectors of the historical SSPs.
-EOF decomposition, also known as principal component analysis, is a statistical method used to analyze structural features in matrix data and extract major data feature quantities.
-Through EOF decomposition, the change information of the original sound velocity field can be easily condensed into the first few principal components and their corresponding spatial functions, so the main information of the sound velocity field can be expressed only by the first few eigenvectors.
+For a specific single coordinate, the configuration of data fusion is presented in figure 2.
+This configuration encompasses real-time SST data, coordinates of latitude and longitude, along with the principal three eigen-vectors of the EOF of historical SSPs.
+EOF decomposition, alternatively referred to as principal component analysis, is a statistical approach employed for examining the structural characteristics within matrix data and extracting the primary quantitative features of the data.
+By means of EOF decomposition, the variation information of the original sound velocity field can be readily compressed into a few principal components and their corresponding spatial functions.
+As a result, the essential information of the sound velocity field can be represented solely by the primary few eigen-vectors.
 To obtain the EOF feature vectors, corresponding 5-year (2015 to 2019) historical SSPs from the target ocean area are first selected and then linearly interpolated with a step size of 1 m.
@@
 […]
-Finally, by fusing data from 8 coordinate points, input training data can be obtained. For location $L_{n-1,m-1}$, the fused data input $\mathcal{F}^{X}_{n-1,m-1}$ can be expressed as:
+Finally, by fusing data from 8 coordinates, input training data can be obtained. For location $L_{n-1,m-1}$, the fused data input $\mathcal{F}^{X}_{n-1,m-1}$ can be expressed as:
 $\mathcal{F}^{X}_{n-1,m-1}=[\boldsymbol{T^{S}_{n-1,m-1}},\boldsymbol{L_{n-1,m-1}},\mathcal{E_{H,K}}],$ (8)
-where $\boldsymbol{T^{S}_{n-1,m-1}}=[T^{s}_{n-1,m-1,1},T^{s}_{n-1,m-1,2},...,T^{s}_{n-1,m-1,H}]^{T}$ is actually a vector composed of H copies of $T^{s}_{n-1,m-1}$ at location $L_{n-1,m-1}$, and $\boldsymbol{L_{n-1,m-1}}$ is similar that $\boldsymbol{L_{n-1,m-1}}=[L_{n-1,m-1,1},L_{n-1,m-1,2},...,L_{n-1,m-1,H}]^{T}$.
+where $\boldsymbol{T^{S}_{n-1,m-1}}=[T^{s}_{n-1,m-1,1},...,T^{s}_{n-1,m-1,H}]^{T}$ is actually a vector composed of H copies of $T^{s}_{n-1,m-1}$ at location $L_{n-1,m-1}$, and $\boldsymbol{L_{n-1,m-1}}$ is similar that $\boldsymbol{L_{n-1,m-1}}=[L_{n-1,m-1,1},L_{n-1,m-1,2},...,L_{n-1,m-1,H}]^{T}$.
 The label output data is $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$ at location $L_{n,m}$.
 III-C SA-MDF-CNN Model
-The network structure of SA-MDF-CNN is shown in Figure 3, mainly including attention module and convolutional network module.
-The first layer is the input layer, the size of which is 1976*6*8, including 6 types of physical quantities (remote sensing SST, latitude, longitude, and the first 3 order eigenvectors) at 8 coordinates with a total depth of 1976 meters.
-The second layer is a flattening layer, which flattens the three-dimensional fusion data into one-dimensional sequence data as input for the self-attention layer.
-The third layer comes the self-attention layer, which converts the fusion data into query, key and value vectors through three linear transforms.
-Then, the dot product of each query vector and each key vector is calculated to obtain the attention score matrix, which is further converted into attention weights through softmax function.
-Finally, the value vectors are summed using attention weights to obtain the final output features.
-These layers are designed to enhance the model’s ability of paying attention to the relationship between SSPs in different ocean areas in the input fusion data, so as to better capture the global dependence and key information between fusion data and the sound velocity distribution in the target ocean area.
+The architectural framework of the SA-MDF-CNN is illustrated in Figure 3, mainly including attention module and convolutional network module.
+The initial layer is the input layer, the size of which is 1976*6*8, including 6 types of physical quantities (remote sensing SST, latitude, longitude, and the first 3 order eigenvectors) at 8 coordinates with a total depth of 1976 meters.
+Then the flattening layer serves as a flattening operation, transmuting three-dimensional fusion data into one-dimensional sequence data as input for the self-attention mechanism.
+The following module is designated as the self-attention layer, which transmutes integrated data into query, key, and value vectors via three distinct linear transformation processes.
+Subsequently, the inner product between every query vector and each key vector is computed, resulting in an attention score matrix.
+This matrix is subsequently transformed into attention weights via the softmax function.
+The final output features are generated through the aggregation of value vectors, facilitated by the application of attention weights.
+The architectural design of these layers is intended to augment the model’s capacity for focusing on the inter - relationships between SSPs situated across disparate oceanic zones within the input fusion data, so as to better capture the global dependence and key information between fusion data and the sound velocity distribution in the target ocean area.
 The fourth and fifth layers are functional layers designed to reshape one-dimensional sequence data into two-dimensional data for subsequent feature extraction through convolution.
-The convolution layer is designed to further extract the features from the output data of the attention layer, which is better at capturing small-scale relationships of input features.
-The pooling layer is used to reduce the spatial dimension of the data, thereby reducing the number of parameters and reduce computational complexity, while retaining important feature information.
-The final two layers are the fully connected layer and the regression layer.
-The fully connected layer integrates the local features into the global features and outputs them to the regression layer.
-The RMSE is designed to be the loss function for calculating the gradient of the model parameters and updating weight parameters of the model through back propagation (BP) algorithm:
+The convolution layer is intended to derive additional features from the resultant data of the attention layer, which is better at capturing small-scale relationships of input features.
+The pooling layer serves to diminish the spatial dimensionality, thereby reducing the number of parameters and reduce computational complexity, while retaining important feature information.
+The penultimate and ultimate layers comprise the fully connected and the regression layer.
+The fully connected layer integrates local features into global representations, subsequently passing them to the regression layer.
+The RMSE is designated as the loss function, which is employed to compute the gradient of the model parameters and update the model’s weight parameters via the back propagation algorithm.
 $Loss=\sqrt{\frac{\sum_{h=1}^{H}(\hat{s}^{Y}_{n,m,h}-s^{Y}_{n,m,h})^{2}}{H}},$ (9)
@@
 Figure 4: The principle of the multi-head self-attention mechanism.
-The self-attention mechanism is a special type of attention mechanism used to process sequential data such as text, images, or time series.
-It allows the model to process the sequence in a way that considers the relationship between each element in the sequence and all other elements.
+The self-attention mechanism constitutes a distinct variant of the attention mechanism, primarily employed for processing temporal data such as text, images, or time-series data.
+It enables the model to analyze the sequence in a manner that takes into account the inter-element relationships within the sequence.
 The self-attention mechanism calculates weights by evaluating the similarities between different elements in a sequence to determine how much an element is related to other elements.
@@
 In order to better capture the associations between different types of SSP, multiple attention heads are used in this paper as shown in figure 4.
-The diagram presented in Figure 4 illustrates the intricate workings of the multi-head self-attention mechanism, which is a fundamental building block in the architecture of transformer models.
-Three pivotal components are involved in the self-attention mechanism: the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ vectors.
-These vectors are derived from the input data through linear transformations, as depicted in the upper right section of the diagram where the input is projected into the query, key, and value spaces using learnable weight matrices.
+The diagram presented in Figure 4 illustrates the intricate workings of the multi-head self-attention mechanism, which is a fundamental building block in the architecture of transformer models, depending on three integral elements: the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ vectors.
+These vectors are derived from the input data through linear transformations, as depicted in the upper right section of the diagram.
 The query vector serves to assess the relevance or correlation with the key vectors.
@@
 In contrast, the value vectors encapsulate the specific features or numeric values associated with each element, playing a crucial role in the aggregation of information.
-The mechanism operates by first computing the dot product between the query and each key vector.
-To prevent the dot products from growing too large and causing numerical instability, especially when the dimensionality of the keys $d_{k}$ is high, each key is normalized by dividing it by $\sqrt{d_{k}}$.
-This normalization step is shown in the diagram as part of the attention calculation process.
-Following this, the softmax function is applied to the normalized dot products to generate a set of attention weights.
-These weights are then used to compute a weighted sum of the value vectors, as represented by the equation:
-$\text{Attention}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}$ (10)
-This operation is performed for each head in the multi-head self-attention mechanism, as shown in the lower section of the diagram.
-Each head attends to different parts of the input sequence, allowing the model to capture a diverse range of dependencies and nuances within the data.
-The outputs from all heads are then concatenated and subjected to a linear transformation, resulting in the final output of the multi-head self-attention mechanism.
-Given an input vector $\mathbf{x}\in\mathbb{R}^{dim}$, represented as $\mathcal{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$, where $dim$ denotes the number of feature dimensions and $M$ represents the number of sequence samples.
-Upon receiving a query vector $\mathbf{q}$, the self-attention function is defined as:
-$\mathrm{att}(\mathcal{X},\mathbf{q})=\sum_{i=1}^{M}a_{i}\mathbf{x}_{i}=\sum_{i=1}^{M}\frac{\exp(\mathrm{s}(\mathbf{k}_{i},\mathbf{q}))}{\sum_{j=1}^{M}\exp(\mathrm{s}(\mathbf{k}_{j},\mathbf{q}))}\mathbf{v}_{i},$ (11)
-where $a_{i}$ is the $i$th attention weight, computed by the softmax function over the scaled dot product of the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$.
-The function $\mathrm{s}(\mathbf{k}_{i},\mathbf{q})$ represents the similarity score between the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$, which is typically calculated as the dot product $\mathbf{k}_{i}^{\top}\mathbf{q}$ divided by the square root of the key vector’s dimension $\sqrt{d_{k}}$.
-The value vector $\mathbf{v}_{i}$ corresponds to the $i$th sample in the input sequence, and the output of the self-attention function is a weighted sum of these value vectors, where the weights are determined by the attention mechanism.
-In a standard self-attention module, the multi-head attention mechanism employs multiple query vectors $\mathcal{Q}$ to select multiple sets of information from the input data in parallel.
+The operation of the mechanism commences with the calculation of the dot-product between the query and every key vector.
+To avoid an excessive increase in the dot-products that could trigger numerical instability, particularly when the dimensionality of the keys $d_{k}$ is substantial, each key is standardized through division by $\sqrt{d_{k}}$.
+This standardization stage is depicted in the diagram as an element of the attention computation procedure.
+Afterward, the softmax function is employed on the standardized dot-products to yield a collection of attention weights.
+Subsequently, these weights are utilized to calculate a weighted sum of the value vectors, as expressed by the equation:
+$\text{Att}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}$ (10)
+This operation is carried out for every head within the attention mechanism, as illustrated in the lower part of the diagram.
+Each head focuses on distinct portions of the input sequence, enabling the model to grasp a wide variety of dependencies and subtleties in the data.
+Subsequently, the outputs of all heads are concatenated and undergo a linear transformation, culminating in the ultimate output of the attention mechanism.
+Given an input vector $\mathbf{x}\in\mathbb{R}^{dim}$, represented as $\mathcal{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$, where $dim$ signifies the number of feature dimensions and $M$ indicates the number of sequence samples.
+Once a query vector $\mathbf{q}$ is received, the self-attention function is formulated as:
+$\mathrm{Sa}(\mathcal{X},\mathbf{q})=\sum_{i=1}^{M}a_{i}\mathbf{x}_{i}=\sum_{i=1}^{M}\frac{\exp(\mathrm{s}(\mathbf{k}_{i},\mathbf{q}))}{\sum_{j=1}^{M}\exp(\mathrm{s}(\mathbf{k}_{j},\mathbf{q}))}\mathbf{v}_{i},$ (11)
+where $a_{i}$ is the $i$th attention weight, computed by the softmax function over the scaled dot-product of the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$.
+The function $\mathrm{s}(\mathbf{k}_{i},\mathbf{q})$ represents the similarity score between the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$, which is typically calculated as the dot-product $\mathbf{k}_{i}^{\top}\mathbf{q}$ divided by the square root of the key vector’s dimension $\sqrt{d_{k}}$.
+The value vector $\mathbf{v}_{i}$ corresponds to the $i$th sample in the input sequence.
+In the self-attention module, the multi-head attention mechanism employs multiple query vectors $\mathcal{Q}$ to select multiple sets of information from the input data in parallel.
 Specifically, each head independently computes the attention weights between the query vectors $\mathcal{Q}$, key vectors $\mathcal{K}$, and value vectors $\mathcal{V}$ to extract different subsets of information.
-The output of the multi-head self-attention module is calculated by concatenating the outputs of all heads and then applying a linear transformation.
+The resultant representation of the multi-head self-attention mechanism is derived through the concatenation of results from each attention head, followed by a linear projection operation.
 The mathematical expression is:
 $\mathrm{MultiHead}(\mathcal{Q},\mathcal{K},\mathcal{V})$ $=\mathrm{Concat}(\mathrm{head}_{1},\ldots,\mathrm{head}_{\mathrm{h}})\mathcal{W}^{O},$ (12)
-where $\mathrm{head_{i}}=\text{Attention}(\mathcal{QW}_{i}^{Q},\mathcal{KW}_{i}^{K},\mathcal{VW}_{i}^{V})$, $\mathcal{W}_{i}^{Q}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{K}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{V}\in\mathbb{R}^{d_{v}}$, and $\mathcal{W}^{o}\in\mathbb{R}^{hd_{v}}$ is a projection of the parameter matrix.
+where $\mathrm{head_{i}}=\text{Att}(\mathcal{QW}_{i}^{Q},\mathcal{KW}_{i}^{K},\mathcal{VW}_{i}^{V})$, $\mathcal{W}_{i}^{Q}\in\mathbb{R}^{d_{k}}$, $\mathcal{W}_{i}^{K}\in\mathbb{R}^{d_{k}}$, $\mathcal{W}_{i}^{V}\in\mathbb{R}^{d_{v}}$, and $\mathcal{W}^{o}\in\mathbb{R}^{hd_{v}}$ is a projection of the parameter matrix.
 IV Results and discussions
-To test the effectiveness of the proposed SA-MDF-CNN model, the monthly average fusion data from July 2015 to 2019 were selected as input training data and the monthly average SSP from 2020 were set to be the output label data.
+To test the effectiveness of the proposed SA-MDF-CNN model, the monthly average fusion data from July 2015 to 2019 were selected as input training data and the monthly average SSPs from 2020 were set to be the output label data.
 Experimental results were compared with other widely used methods for constructing spatial sound velocity distribution, including convolutional neural network (CNN), spatial interpolation (SITP) and mean values (MEAN), respectively.
@@
 maximum epochs 100
-learning rate $10^{-3}$
+learning rate 0.001
 drop factor of learning rate 0.1
@@
 Figure 5: Remote sensing SST data.
-To test the accuracy performance of proposed model in estimating SSPs, real-time remote sensing SST data from different locations and depths were fused with EOF feature vectors to construct the sound velocity field.
-The remote sensing SST data are shown in figure 5, and a visual display of the fusion data is given in figure 6.
+To generate the fusion data necessary for evaluating the accuracy of the recommended model in estimating the sound velocity field, real-time remote sensing SST data from multiple locations were combined with the EOF feature vector to reconstruct the sound velocity field.
+Figure 5 shows the remote sensing SST data, and figure 6 presents a visual depiction of the fused data.
 Figure 6: Fusion data of remote sensing SST data, latitude and longitude coordinates, and EOF feature vectors.
-Figure 7: Comparison of real-time estimated SSP results of different algorithms with depth of 1976 m, where (a) is the SA-MDF-CNN algorithm, (b) is the CNN algorithm, (c) is the spatial interpolation algorithm, and (d) is the mean value method.
+Figure 7: Evaluation of real-time SSP estimation outcomes across diverse algorithmic approaches at 1976 meter bathymetric measurement, where (a) is the SA-MDF-CNN algorithm, (b) is the CNN algorithm, (c) is the spatial interpolation algorithm, and (d) is the mean value method.
 TABLE II: Comparison of RMSE at the same depth:1976 meters.
@@
 Average 0.1468 0.1661 0.2326 0.2447
+Figure 8: A comparison example of real-time estimated SSP results of different algorithms with depths ranging from 0 to 1976 meters, where (a) is the comparison of estimated results and real SSP, and (b) is the comparison of error distribution between the estimated results of different algorithms and the real SSP.
 The real-time SSP estimation results of SA-MDF-CNN algorithm with depths ranging from 0 to 1976 meters within $7^{\circ}N$ to $28^{\circ}N$, $150^{\circ}E$ to $165^{\circ}E$ are compared with those of CNN, SITP and mean values method as shown in figure 7.
-From the intuitive representation of the sound velocity heat-map, it can be seen that the sound velocity distribution estimated by SA-MDF-CNN and CNN has a significant improvement in accuracy compared to SITP and Mean value methods, and the estimated values of the proposed model are more similar to the sound velocity distribution of real samples.
+From the intuitive representation of the sound velocity heat-map, the sound velocity distribution estimated by the SA-MDF-CNN and CNN models exhibits a substantial enhancement in accuracy relative to the SITP and Mean value methods as evidenced by the results, and the estimated values of the proposed model are more similar to the sound velocity distribution of real samples.
 A comprehensive evaluation of the accuracy performance of various algorithms across different locations at depths ranging from 0 to 1976 meters is presented in Table II.
@@
 However, among all methods the error disturbance of SA-MDF-CNN is the smallest.
-Figure 8: A comparison example of real-time estimated SSP results of different algorithms with depths ranging from 0 to 1976 meters, where (a) is the comparison of estimated results and real SSP, and (b) is the comparison of error distribution between the estimated results of different algorithms and the real SSP.
 Figure 9: Comparison of estimation results of different algorithms at different shallow ocean depths.
@@
 Figure 10: Comparison of error distributions of real-time estimated SSPs for different algorithms at different shallow sea depths.
-To further test the accuracy and performance of the model in estimating the sound velocity distribution in shallow water areas, the real-time SSP estimation results of SA-MDF-CNN algorithm at 200 m, 300 m and 500 m depths are compared with those of CNN, SITP and mean values method as shown in figure 9.
-As can be seen from the enlarged part of each sub-figure, the estimated SSP curves by SA-MDF-CNN at 200 m, 300 m and 500 m depths are significantly closer to the real SSP prediction results than those of CNN, SITP and mean method at the same ocean area.
-Specifically, an in-depth investigation into the performance of different algorithms in shallow ocean areas at varying depths and locations was conducted, with the real-time SSP estimation RMSE results compared and displayed in Table LABEL:table3.
+To rigorously evaluate the precision and performance of the model for reconstructing acoustic velocity profiles in coastal marine environments, a comparative analysis was conducted between the instantaneous SSP reconstructions generated by the SA-MDF-CNN framework at 200 m, 300 m, and 500 m depth layers and those produced by conventional CNN architectures, the SITP technique, and statistical averaging approaches, as illustrated in Figure 9.
+The magnified sections of individual subplots reveal that SA-MDF-CNN-derived SSP estimations at 200 m, 300 m, and 500 m depth layers demonstrate enhanced alignment with ground-truth SSP predictions when compared to conventional CNN architectures, SITP algorithms, and mean method within identical marine regions.
+Specifically, an in-depth investigation into the efficacy of various algorithms within shallow ocean areas at varying depths and locations was conducted, with the real-time SSP estimation RMSE results compared and displayed in Table LABEL:table3.
 The largest RMSE is observed at a depth of 200 meters, which can be attributed to the increased interference from real-time sea conditions in the shallow sea environment, such as wind speed, wave height, tsunami, and other factors.
@@
 These experimental results collectively illustrate that the proposed model maintains significant precision advantages even in shallow sea environments of varying depths, thereby underscoring its applicability and effectiveness.
-The results indicate that SA-MDF-CNN is also applicable even in shallow ocean environment with different depths, and still has smaller RMSE compared with other algorithms, which means that the model still has good feature capture ability in shallow ocean scenes with significant changes in sound velocity.
+Experimental findings reveal the SA-MDF-CNN architecture maintains operational efficacy across heterogeneous depth profiles within coastal marine conditions, while exhibiting RMSE relative to conventional methodologies, which means that the model still has good feature capture ability in shallow ocean scenes with significant changes in sound velocity.
 The average absolute error distribution comparison of real-time estimated SSP of different algorithms tested in the set of 200m and 500m depths is given in Figure 10.
@@
 To investigate why the proposed SA-MDF-CNN model has better performance compared to CNN, we conducted an experiment and visually represented the attention weights of the model as shown in figure 11.
-Figure 11 (a) illustrates the first three order EOF feature vectors of the historical SSPs at various latitude and longitude coordinates.
+Figure 11 (a) illustrates the primary three-order EOF eigenvectors corresponding to historical SSPs at multiple geospatial coordinates.
 Figure 11 (b), (c), and (d) correspond to the attention weights when the training epochs are 10, 50, and 100, respectively.
-Through observing the evolution trend of attention focus with increasing training times, we found that the SA-MDF-CNN model increasingly focuses on the sound velocity distribution in shallow waters within 1000 meters.
-This trend aligns with the first three orders of EOF feature distribution shown in panel (a), although the EOF itself emphasizes the SSP distribution between 0-500 meters.
+Through observing the evolution trend of attention focus with increasing training times, we found that the SA-MDF-CNN model increasingly focuses on the acoustic velocity profile in shallow marine environments within 1000 meters.
+This trend aligns with the first three orders of EOF feature distribution shown in figure 11 (a), although the EOF itself emphasizes the SSP distribution between 0-500 meters.
 The attention depth of the proposed model spans 0-1500 meters, which is attributed to the variation of SSP distribution under different latitude and longitude coordinates.
-Therefore, in order to enhance the model’s ability to focus on the global SSP distribution, we added a multi-head self-attention layer after input data fusion.
-This mechanism further confirms the effectiveness of the proposed model.
+To improve the capability of capturing global SSP distribution patterns, a multi-head self-attention mechanism was incorporated following the fusion of input data.
+This architectural enhancement demonstrates the robustness and validity of the introduced framework.
 Experiments based on the aforementioned public dataset demonstrate that the proposed SA-MDF-CNN model outperforms other algorithms in estimating sound velocity in both shallow and deep waters.
@@
 Although the accuracy performance in SSP estimation of SA-MDF-CNN is better than traditional methods, the disadvantage is that SA-MDF-CNN increases the number of parameters and computational overhead.
-Table IV shows the comparison of parameters and running time between the proposed model and CNN.
-It can be seen that the proposed model in this paper needs more parameters and training time than CNN, which is a necessary overhead for performance improvement.
+Table IV presents the computational complexity analysis and temporal efficiency comparison between SA-MDF-CNN and conventional CNN architectures.
+The experimental results indicate that the introduced methodology requires increased parameter quantities and extended training durations compared to CNN, representing an essential computational investment for achieving enhanced predictive capabilities.
 Figure 12: Comparison of RMSE convergence of different algorithms.
@@
 Figure 13: Data sampling locations of ocean experiments.
-Figure 14: Comparison of real-time SSP estimation results of different algorithms on ocean experimental data, where (a) shows the SSP curves, and (b) shows the error distributions.
-In order to verify the practicality of the SA-MDF-CNN model, a comprehensive deep-ocean experiment was conducted in the South China Sea in April 2023.
-As illustrated in figure 13 (a), various instruments such as CTD and XCTD were employed at diverse marine locations, yielding a total of 14 distinct samples.
+In order to verify the practicality of the SA-MDF-CNN model,a deep-sea experiment was conducted in the South China Sea in April 2023.
+As illustrated in figure 13, various instruments such as CTD and XCTD were employed at diverse marine locations, yielding a total of 14 distinct samples.
 The depth of sampling reached 3500 meters, with intervals set at every 100 meters.
 It is evident that the selected location for this experiment does not conform to the rigid 3x3 grid pattern, thus allowing for a more rigorous evaluation of the proposed model.
-The fusion data at the blue coordinates in figure 13 (b) were selected as the input of the model, and the real-time SSP under the red pentagram was used as the prediction output of the model.
-Figure 15: Comparison of real-time estimation SSP results of different algorithms in shallow sea environment with different depths in ocean experiment, where (a) and (c) are SSP curves of 300 meters depth, (b) and (d) are error distributions of 500 meters depth.
-Figure 16: RMSE comparison of real-time estimated SSPs by different algorithms of ocean experiments.
+The input data for the model were extracted from the fusion dataset corresponding to the blue coordinates, while the real-time SSP associated with the red pentagram served as the model’s predictive output.
+Figure 14: Evaluation of real-time SSP estimation outcomes utilizing diverse algorithms on oceanographic experimental data, where (a) shows the SSP curves, and (b) shows the error distributions.
+Figure 15: Comparison of real-time estimation SSP results of different algorithms in shallow sea environment with different depths in ocean experiment, where (a) and (b) are SSP curves of 300 meters depth, (c) and (d) are error distributions of 500 meters depth.
 To evaluate the effectiveness of the model in estimating the distribution of sound velocity at full-ocean depth, the estimated SSP of SA-MDF-CNN is compared with other algorithms in figure 14.
@@
 To further test the sound velocity estimation performance in shallow waters with significant changes in sound velocity, estimation of sound velocity at 300, 500 depth meters by SA-MDF-CNN are compared with other methods in figure 15.
-The results indicate that the proposed algorithm has smaller error disturbances and maintains significant accuracy advantages compared to other algorithms in shallow water environments.
+The experimental outcomes demonstrate that the introduced methodology exhibits reduced error fluctuations and sustains notable precision benefits relative to alternative approaches in shallow marine settings.
+This comparative analysis highlights the robustness and superior performance of the proposed technique under such environmental conditions.
 Figure 16 gives a more detailed comparison of the RMSE results of different algorithms, clearly showing that the proposed model has significant advantages at different depths.
+Figure 16: RMSE comparison of real-time estimated SSPs by different algorithms of ocean experiments.
 VI Conclusion
-To construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, a SA-MDF-CNN model driven by multimodal data fusion is proposed in this paper, which fuses historical SSP features, real-time remote sensing SST, latitude and longitude coordinates.
-The fusion data is constructed by gridding, and the real-time SSP is retrieved based on the real-time remote sensing SST.
-The effectiveness of our proposed model was tested in the case of non-standardized grid through ocean experiments.
-The experimental results show that our proposed model is not only suitable for real-time SSP estimation under non-strictly standardized grid conditions, but also suitable for shallow sea conditions with different depths.
+To construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, an SA-MDF-CNN model driven by multimodal data fusion is proposed in this paper, which fuses historical SSP features, real-time remote sensing SST, latitude and longitude coordinates.
+The attention mechanisms were applied to enhance the model’s capacity to learn complex interactions within multimodal data.
+To evaluate the accuracy performance of the model, we conducted experiments using Argo data from the Pacific region and further validated it through non-standardized grid ocean experiments in the South China Sea in April 2023.
+The empirical findings demonstrate that the developed framework exhibits dual applicability, functioning effectively for real-time SSP reconstruction in non-uniform grid environments while maintaining effectiveness across diverse shallow-water and deep-water scenarios.
+At the same time, in order to enhance the interpretability of the model, we visualized the model’s parameters and found that the parameter weights were more concentrated in shallow water parts, which is consistent with the magnitude of the change in sound velocity.
+SA-MDF-CNN provides a fast and accurate way to estimate the distribution of regional sound velocity, which is suitable for areas with stable marine environments.
+However, it still has some limitations, for example, the model is insensitive to fluctuations in sound velocity distribution caused by internal ocean waves.
+Therefore, in future work, we will focus on studying the accurate estimation of sound velocity distribution under the influence of ocean internal waves without conducting on-site measurements of underwater data.
 Acknowledgments
@@
 The remote sensing SST data can be acquired at the National Oceanic and Atmospheric Administration (https://www.commerce.gov/, latest access: December 25, 2024).
+There is no conflict of interest.